OpenStudy (anonymous):

how do i simplify 5n^2+3n well how do i simplify quadratics with no "C" is it the same as one with a "C"

OpenStudy (anonymous):

$$5n^2+3n=n(5n+3)=5n(n+3/5)$$ Is this the form you need?

OpenStudy (anonymous):

You can simplify it at the same way as any quadratic,only keeping in mind that C equals 0.

OpenStudy (anonymous):

i think the way its written the second time is a good enough anwser thanks :)